OpenStudy (anonymous):

1. Choose the correct x-intercept and y-intercept of the equation 2x − y = 4. x-intercept: (−4, 0), y-intercept: (0, 2) x-intercept: (−4, 0), y-intercept: (2, 0) x-intercept: (2, 0), y-intercept: (0, −4) x-intercept: (0, 2), y-intercept: (0, −4) 2. Choose the correct x-intercept and y-intercept of the equation 3x − 2y = 6. x-intercept: (0, −3), y-intercept: (0, 2) x-intercept: (−3, 0), y-intercept: (2, 0) x-intercept: (0, 2), y-intercept: (0, −3) x-intercept: (2, 0), y-intercept: (0, −3) 3. Choose the graph that corresponds to the equation 3x + 2y = 12.

OpenStudy (texaschic101):

To find the x intercept, sub in 0 for y To find the y intercept, sub in 0 for x 2x - y = 4 2x - 0 = 4 2x = 4 x = 2 --- x intercept is (2,0) 2(0) - y = 4 -y = 4 y = -4 --- y intercept is (0,-4) Its that simple...now you can do the next one :)